### v1

@@ Abstract
Large Multimodal Models (LMMs) have demonstrated remarkable capabilities. While existing benchmarks for evaluating LMMs mainly focus on image comprehension, few works evaluate them from the image generation perspective. To address this issue, we propose a straightforward automated evaluation pipeline. Specifically, this pipeline requires LMMs to generate an image-prompt from a given input image. Subsequently, it employs text-to-image generative models to create a new image based on these generated prompts. Finally, we evaluate the performance of LMMs by comparing the original image with the generated one. Furthermore, we introduce MMGenBench-Test, a comprehensive benchmark developed to evaluate LMMs across $13$ distinct image patterns, and MMGenBench-Domain, targeting the performance evaluation of LMMs within the generative image domain. A thorough evaluation involving over 50 popular LMMs demonstrates the effectiveness and reliability in both the pipeline and benchmark. Our observations indicate that numerous LMMs excelling in existing benchmarks fail to adequately complete the basic tasks, related to image understanding and description. This finding highlights the substantial potential for performance improvement in current LMMs and suggests avenues for future model optimization. Concurrently, our pipeline facilitates the efficient assessment of LMMs performance across diverse domains by using solely image inputs.
Figure 1: The MMGenBench-Test Benchmark consists of $13$ image patterns, each of which includes several images. The text, accompanied by a corresponding pattern, serves as a concise explanation of that specific image pattern. Please refer to the Appendix B.1 for more details.
@@ 1 Introduction
We have witnessed the quick progress of large multimodal models (LMMs) [7, 5, 28, 50, 21, 39, 12], which efficiently utilize the strength of LLMs [34, 37, 33, 40, 23, 32] in processing both visual and textual inputs. Compared to text, visual images are characterized by their high level of abstraction and information density, while exhibiting strong spatial correlations and structural complexity. The development of comprehensive benchmarks is crucial for enhancing and accurately evaluating LMMs. Specifically, many popular benchmarks [27, 22, 45, 42, 49, 16, 17, 29], provide standardized evaluations for LMMs by assessing multimodal tasks across various datasets. However, these benchmarks frequently depend on traditional datasets, resulting in issues of data leakage and limited task diversity. Although many various tasks such as VQA [4], Image Caption [24], and OCR [46] are exquisitely designed, most existing benchmarks only focus on evaluating the understanding performance of LMMs.
@@
On the other hand, the understanding and generation of images remain disparate fields, with the most formidable models [15, 1, 33] in their respective domains adhering to distinct paradigms. For instance, GPT-4 [33], which is grounded in the next token prediction paradigm, exhibits a formidable capacity for image comprehension. Similarly, Flux [1] has achieved noteworthy success in text-to-image synthesis by leveraging diffusion models [20, 35]. This divergence underscores the complexity of achieving a unified approach to image processing and synthesis, as the state-of-the-art techniques continue to evolve along separate trajectories. Furthermore, Large Multimodal Models (LMMs) are extensively employed to generate training data for generative models. It is noteworthy that LMMs excel in image-to-text tasks, while diffusion models are particularly effective in text-to-image tasks. A robust understanding of an image implies that LMMs can distill its essential information into text prompts, which can then be used by text-to-image models to reconstruct the scene to a certain extent. This process can be viewed as a form of “compression”.
Hence, it is both reasonable and significant to evaluate the performance of LMMs using diffusion models. Our work aims to bridge this gap by providing a comprehensive evaluation pipeline.
@@
In this paper, we propose MMGenBench-Test, a comprehensive benchmark designed to evaluate LMMs across 13 distinct image patterns (refer to Fig. 1), and MMGenBench-Domain, which focuses on assessing LMMs performance within the generative image domain. To achieve this, we introduce a pipeline that initially allows LMMs to generate image-prompt from input images, then employs text-to-image generative models to create new images. Finally, we use an image representation model to obtain the embeddings of images and perform post-processing to assess the performance of LMMs in image understanding and description. We conduct extensive experiments on over 50 popular LMMs, demonstrating the effectiveness and reliability of both the pipeline and benchmark. Notably, our findings reveal that numerous LMMs excelling in existing benchmarks fail to address the basic tasks of image understanding and description, highlighting the substantial potential for improvement and optimization in future models. This straightforward automated evaluation pipeline also enables efficient assessment of LMMs across diverse domains through image inputs alone, providing a flexible and scalable benchmarking tool.
@@
  • The proposed MMGenBench is the first automated pipeline, designed to evaluate the capabilities of LMMs in image understanding and description by solely utilizing images. This pipeline utilizes text-to-image models and image representation models for automated evaluation, thereby markedly minimizing human involvement and improving the efficiency and objectivity of the evaluation procedure.
@@
  • We developed MMGenBench-Test, a comprehensive benchmark designed to evaluate LMMs across $13$ image patterns, and MMGenBench-Domain, which assesses the performance of LMMs in the generative image domain.
@@
  • Our study includes a broad evaluation of over 50 popular LMMs, providing critical insights into their capabilities and limitations in basic image understanding and description tasks.
Please refer to the Appendix A for related work.
@@
Figure 2: An overview of the MMGenBench pipeline, illustrating the automated evaluation process. It starts by receiving a user input (including the task instruction prompt and input images), and then generates the corresponding image-prompt. Subsequently, this process is followed by using a powerful text-to-image model to generate new images, then produces the representation of the input images and the generated ones using an image representation model, and finally outputs the evaluation score of LMMs.
2 The MMGenBench Benchmark
2.1 Evaluation Pipeline
@@
The proposed pipeline for LMMs, based on text-to-image generative models and image representation models, consists of three components: image-prompt generation, new image generation, and quantitative metric computation.
Image-prompt Generation.
@@
The input modalities for the LMM primarily consist of two parts: the image used for model understanding and inference, and the prompt guiding the inference direction. To facilitate a standardized workflow, we employ a manually crafted, normalized prompt $\mathbf{P}_{art}$. This prompt constrains the LMM’s comprehension and reasoning across five dimensions: role, definition explanation, task instruction, key points and requirements, and output format (see Appendix C.1 for details). This process can be formalized as follows:
$\mathbf{P}_{gen}=\text{LMM}(\mathbf{I},\mathbf{P}_{art}).$ (1)
@@
Through the above process, we obtain the image-prompt (the detailed description of an image) $\mathbf{P}_{gen}$ output by the LMM, which will serve as the input for the subsequent stage.
New Image Generation.
@@
In this stage, the main process involves the utilization of text-to-image models to generate new images based on the given image-prompt. Theoretically, the generation quality is highly dependent on the chosen text-to-image model. To minimize variable interference, we standardize the evaluation by selecting four state-of-the-art models: FLUX.1-dev [1], Stable Diffusion 3.5 [15], Kolors [2] and Lumina [18]. The comparison results across these models provide cross-validation of generation effectiveness. This phase is represented as:
@@
Quantitative Metric Computation.
We quantify the functionality of LMM by evaluating the similarity between the input image $\mathbf{I}$ and the generated one $\mathbf{I}_{gen}$. Since most generative models introduce a certain degree of randomness in their inference process, achieving pixel-level consistency between images is challenging. To address this, we conduct comparisons at the representational level. Specifically, we utilize the Unicom model [6] to encode each image and obtain its respective representation. This phase is encapsulated as:
$\mathbf{F}_{I}=\text{Encoder}(\mathbf{I}),\mathbf{F}_{I_{gen}}=\text{Encoder}(\mathbf{I}_{gen}),$ (3)
where $\mathbf{F}_{I}$ and $\mathbf{F}_{I_{gen}}$ represent the features extracted from the input image $\mathbf{I}$ and the generated image $\mathbf{I}_{gen}$, respectively. To provide a quantitative evaluation, we compute both a similarity score and a generation quality score based on these feature representations.
@@ 2.2 Evaluation Metric
where $\mu_{x}$ and $\mu_{y}$ denote the means of the input image features $F_{I}$ and the generated image features $F_{I_{gen}}$, respectively, while $\Sigma_{x}$ and $\Sigma_{y}$ represent their covariances. The notation $\text{Tr}(\cdot)$ denotes the trace of a matrix.
2.3 MMGenBench Benchmark Construction
2.3.1 Overview
@@
To effectively measure the understanding and description capabilities of LMMs across various types of images, we constructed a high-quality test set MMGenBench-Test for $13$ image patterns using the JourneyDB test set [36]. We proposed a multi-stage method for extracting and annotating image patterns as illustrated in Fig. 4. To ensure the results’ accuracy, we manually double-checked the image patterns and performed the final annotations.
In addition, we have also constructed a dataset in the “image generation” domain, termed MMGenBench-Domain, to evaluate the ability of LMMs in the understanding and describing “generated images” task. It is important to emphasize that our proposed pipeline can measure the ability of LMMs to understand and describe images in any domain. By utilizing images from a particular domain, we can easily assess the performance of LMMs specific to that domain.
Figure 3: Statistics of MMGenBench-Test, which contains $13$ image patterns with $1,284$ images. More details are in Sec. 2.3.
2.3.2 Dataset Statistics
@@
To construct MMGenBench-Domain, we randomly sampled $10,000$ images from the JourneyDB validation set. By utilizing the proposed pipeline, we can evaluate the image understanding and descriptive performance of LMMs within this domain, obviating the need for additional data.
@@ 2.3.3 Data Collection
To construct MMGenBench-Test, we first extract all images from the JourneyDB test set and process them using the pipeline, shown in Fig. 4. For the domain-specific dataset, we use the JourneyDB validation set for its construction. Subsequently, we will elaborate on each processing step.
Image Pattern Extraction.
@@
We extract image patterns from existing images to measure the understanding and description ability of LMMs across various image categories. Given that we possess only images but no additional information, we utilize the powerful GPT-4o model to extract and analyze the underlying image patterns. The task is executed by GPT-4o in three sequential steps: 1) providing a detailed description of the image; 2) annotating the possible patterns based on the image features and description; and 3) explaining the rationale behind the annotated image patterns. By utilizing our carefully designed prompts (refer to Appendix B.3), GPT-4o can perform the task effectively. An example is presented in Fig. 8 in Appendix B.2, wherein the image description, alongside the extracted image patterns and rationales, aligns accurately with the image content. By annotating each image in the JourneyDB test set, we ultimately obtained a total of $1868$ image patterns.
Image Pattern Summary.
@@
By leveraging GPT-4o, we extracted an extensive range of image patterns and observed that the semantic meanings across different patterns may be consistent. We conducted thorough statistics of the image patterns, by quantifying the occurrences of each type and ranking them in descending order (as seen in Appendix B.1). Subsequently, GPT-4-turbo was utilized to generate the summary. To ensure the accuracy and validity of the extracted patterns, we carefully designed model prompts for this task (shown in Appendix B.3). This procedure ultimately resulted in a hand-crafted list of $13$ distinct image patterns, each thoroughly described in Appendix B.1.
Image Pattern Re-annotation.
@@
We conducted a re-annotation of the JourneyDB test set using the previously identified $13$ image patterns. Similar to Image Pattern Extraction, GPT-4o was employed to generate descriptions for each image. Subsequently, these descriptions were annotated based on the predefined image patterns, accompanied by the underlying reasons. A key difference in this process is that re-annotation is restricted to the specified $13$ image patterns, excluding any other patterns. Given the extensive number of images in the test set, it is essential to select a subset of the images for subsequent operations. The re-annotated images were systematically classified into distinct image patterns, each containing a specified number of images. To construct the final dataset, we randomly sampled a subset of images from each pattern with a selection probability of $\frac{100}{N}$, where $N$ represents the total number of images within the pattern. To mitigate potential annotation errors, we manually verified and annotated each image. The final dataset was established by voting between the model predictions and manual annotations.
Domain Data Collection.
@@
Constructing high-quality datasets is an extremely expensive task. Therefore, it is crucial to develop an efficient and convenient method for creating domain-specific datasets when evaluating the image understanding and description capabilities of LMMs in a new domain. By leveraging the proposed pipeline, our method enables a seamless evaluation of LMM performance across various domains, only depending on the availability of domain-specific images. In this study, we randomly select $10,000$ images from the JourneyDB validation set to create our domain-specific dataset. To quantify LMM performance within this domain, we calculate both FID-Score and SIM-Score.

### v2

@@ Abstract
Large Multimodal Models (LMMs) demonstrate impressive capabilities. However, current benchmarks predominantly focus on image comprehension in specific domains, and these benchmarks are labor-intensive to construct. Moreover, their answers tend to be brief, making it difficult to assess the ability of LMMs to generate detailed descriptions of images. To address these limitations, we propose the MMGenBench-Pipeline, a straightforward and fully automated evaluation pipeline. This involves generating textual descriptions from input images, using these descriptions to create auxiliary images via text-to-image generative models, and then comparing the original and generated images. Furthermore, to ensure the effectiveness of MMGenBench-Pipeline, we design MMGenBench-Test, evaluating LMMs across 13 distinct image patterns, and MMGenBench-Domain, focusing on generative image performance. A thorough evaluation involving over 50 popular LMMs demonstrates the effectiveness and reliability of both the pipeline and benchmark. Our observations indicate that numerous LMMs excelling in existing benchmarks fail to adequately complete the basic tasks related to image understanding and description. This finding highlights the substantial potential for performance improvement in current LMMs and suggests avenues for future model optimization. Concurrently, MMGenBench-Pipeline can efficiently assess the performance of LMMs across diverse domains using only image inputs.
{strip}[Uncaptioned image]
Figure 1: The MMGenBench-Test consists of $13$ distinct image patterns, each of which includes several images. The text, accompanied by a corresponding pattern, serves as a concise explanation of that specific image pattern. Please refer to the Appendix B.1 for more details.
@@ 1 Introduction
We have witnessed rapid progress of Large Multimodal Models (LMMs) [42, 81, 27, 66, 50, 4, 6, 12, 65, 10, 53], which efficiently utilize the strength of LLMs [54, 61, 53, 68, 32, 52] in processing visual and textual inputs. Compared to text, visual images are characterized by their high level of abstraction and information density, while exhibiting strong spatial correlations and structural complexity. The development of comprehensive benchmarks is crucial for enhancing and accurately evaluating LMMs. Specifically, many popular benchmarks [41, 30, 76, 72, 80, 17, 19, 45, 55], provide standardized evaluations for LMMs by assessing multimodal tasks across various datasets. However, these benchmarks frequently depend on traditional datasets, resulting in issues of data leakage and limited task diversity. Although many various tasks such as VQA [2, 49, 79, 51, 47, 24, 55], Image Caption [33, 44], and OCR [77, 45, 18, 74] are carefully designed, as shown in Fig. 2(a), most existing benchmarks require expensive manual construction and focus on specific domains, making it difficult to extend to other domains. The constructed sample answers are mostly brief, focusing only on evaluating the understanding performance of LMMs and ignoring the evaluation of the ability to generate detailed descriptions of images. In addition, the understanding and generation of images remain disparate fields, with the most powerful models [16, 34, 53] in their respective domains adhering to distinct paradigms. For instance, GPT-4 [53], which is grounded in the next token prediction paradigm, exhibits an impressive capacity for image comprehension. Similarly, Flux [34] has achieved noteworthy success in text-to-image synthesis by leveraging diffusion models [25, 58]. This divergence highlights the complexity of achieving a unified approach to image processing and synthesis, as the state-of-the-art techniques continue to evolve along separate trajectories. Furthermore, LMMs are extensively employed to generate data for generative models [56, 70, 39, 28]. It is noteworthy that LMMs excel in image-to-text tasks, while diffusion models are particularly effective in text-to-image tasks. A robust understanding of an image implies that LMMs can distill its essential information into text prompts, which text-to-image models can use to reconstruct the scene to a certain extent. This process can be viewed as a form of “compression”. Hence, it is both reasonable and significant to evaluate the performance of LMMs using diffusion models.
@@
In this paper, we propose MMGenBench-Pipeline, a fully automated evaluation pipeline (refer to Fig. 2(b)) that initially allows LMMs to generate textual descriptions from input images, then employs text-to-image generative models to create auxiliary images. Finally, we use an image representation model to obtain the embeddings of images and perform post-processing to assess the performance of LMMs in image understanding and description. To verify the effectiveness of MMGenBench-Pipeline, we introduce MMGenBench-Test, a comprehensive benchmark designed to evaluate LMMs across 13 distinct image patterns (refer to Fig. 1), and MMGenBench-Domain, which focuses on assessing LMMs performance within the generative image domain. Extensive experiments on over 50 popular LMMs demonstrate the effectiveness and reliability of both the pipeline and benchmark. Notably, our findings reveal that numerous LMMs excelling in existing benchmarks fail to address the basic tasks of image understanding and description, highlighting the substantial potential for improvement and optimization in future models. The MMGenBench-Pipeline also enables efficient evaluation of LMMs across diverse domains through image inputs alone, providing a flexible and scalable benchmarking tool.
@@
  • Fully Automated Evaluation Pipeline: We propose the first fully automated pipeline MMGenBench-Pipeline designed to evaluate the capabilities of LMMs in image understanding and description by solely utilizing images. This pipeline utilizes text-to-image models and image representation models for automated evaluation, thereby markedly minimizing human involvement and improving the efficiency and objectivity of the evaluation procedure.
@@
  • Comprehensive Benchmarks: In order to verify the effectiveness of MMGenBench-Pipeline, we developed MMGenBench-Test, a comprehensive benchmark designed to evaluate LMMs across $13$ image patterns, and MMGenBench-Domain, which assesses the performance of LMMs in the generative image domain.
@@
  • Extensive Evaluation: Our study includes a broad evaluation of over 50 popular LMMs, providing critical insights into their capabilities and limitations in basic image understanding and description tasks.
Please refer to Appendix A for related work.
Figure 2: Comparison between previous benchmarks and MMGenBench. MMGenBench has several novel features: 1) Based on powerful text-to-image models and image representation models, MMGenBench can fully automatically complete the evaluation of LMMs without the need for expensive manual annotation; 2) MMGenBench can easily evaluate the performance of LMMs in any domain, whereas previous benchmarks could mostly only evaluate the performance in specific domains; 3) The “answer” to previous benchmarks were mostly brief, overlooking the basic ability to generate detailed descriptions of images.
@@
Figure 3: An overview of the MMGenBench-pipeline, illustrating the fully automated evaluation process. It starts by receiving user input (including the task instruction prompt and input images), and then generates the corresponding textual descriptions of input images. Subsequently, this process is followed by using a powerful text-to-image model to generate auxiliary images, then produces the representation of the input images and the generated ones using an image representation model, and finally outputs the evaluation score of LMMs.
2 The MMGenBench-Pipeline
2.1 Fully Automated Evaluation Pipeline
@@
The proposed pipeline for LMMs, based on text-to-image generative models and image representation models, consists of four components: test set construction, textual description generation, auxiliary image generation, and quantitative metric computation.
Test Set Construction. Our method can easily evaluate image understanding and description capabilities of LMMs in any domain. As shown in Stage 1 of Fig. 3, given any image from the domain to be tested, we can construct the multi-modal input for the LMMs to be evaluated simply using the predefined prompts, and apply it to the subsequent process.
@@
Textual Description Generation. The input modalities for the LMM primarily consist of two parts: the image used for model understanding and inference, and the prompt guiding the inference direction. To facilitate a standardized workflow, we employ a manually crafted, normalized prompt $\mathbf{P}_{art}$. This prompt constrains the LMM’s comprehension and reasoning across five dimensions: role, definition explanation, task instruction, key points and requirements, and output format (see Appendix C.1 for details). This process can be formalized as follows:
$\mathbf{P}_{gen}=\text{LMM}(\mathbf{I}_{inp},\mathbf{P}_{art}).$ (1)
@@
Through the above process, we obtain the detailed textual description of an image $\mathbf{P}_{gen}$, generated by the LMM, which will serve as the input for the subsequent stage.
@@
Auxiliary Image Generation. In this stage, the main process involves the utilization of text-to-image models to generate auxiliary images based on the given textual descriptions. Theoretically, the generation quality is highly dependent on the chosen text-to-image model. To minimize variable interference, we standardize the evaluation by selecting four state-of-the-art models: FLUX.1-dev [34], Stable Diffusion 3.5 [16], Kolors [62] and Lumina [20]. The comparison results across these models provide cross-validation of generation effectiveness. This phase is represented as:
@@
Quantitative Metric Computation. We quantify the functionality of LMM by evaluating the similarity between the input image $\mathbf{I}_{inp}$ and the generated one $\mathbf{I}_{gen}$. Since most generative models introduce a certain degree of randomness in their inference process, achieving pixel-level consistency between images is challenging. To address this, we conduct comparisons at the representational level. Specifically, we utilize the Unicom [5] model to encode each image and obtain its representations. This phase is encapsulated as:
$\mathbf{F}_{I_{inp}}=\text{Encoder}(\mathbf{I}_{inp}),\mathbf{F}_{I_{gen}}=\text{Encoder}(\mathbf{I}_{gen}),$ (3)
where $\mathbf{F}_{I_{inp}}$ and $\mathbf{F}_{I_{gen}}$ represent the features extracted from the input image $\mathbf{I}_{inp}$ and the generated image $\mathbf{I}_{gen}$, respectively. To provide a quantitative evaluation, we compute both a similarity score and a generation quality score based on these feature representations.
@@ 2.2 Evaluation Metric
where $\mu_{x}$ and $\mu_{y}$ denote the means of the input image features $F_{I_{inp}}$ and the generated image features $F_{I_{gen}}$, respectively, while $\Sigma_{x}$ and $\Sigma_{y}$ represent their covariances. The notation $\text{Tr}(\cdot)$ denotes the trace of a matrix.
3 The MMGenBench Benchmark
3.1 Overview
@@
To effectively measure the understanding and description capabilities of LMMs across various types of images, we constructed a high-quality test set MMGenBench-Test for $13$ image patterns using the JourneyDB [60] test set. We proposed a multi-stage method for extracting and annotating image patterns as illustrated in Fig. 5. To ensure the results’ accuracy, we manually double-checked the image patterns and performed the final annotations.
In addition, we have also constructed a dataset in the “image generation” domain, termed MMGenBench-Domain, to evaluate the ability of LMMs in the understanding and describing “generated images” task. It is important to emphasize that MMGenBench-Pipeline can measure the ability of LMMs to understand and generate detailed image descriptions in any domain. By utilizing images from a particular domain, we can easily assess the performance of LMMs specific to that domain.
Figure 4: Statistics of MMGenBench-Test, which contains $13$ image patterns with $1,284$ images. More details are in Sec. 3.
3.2 Dataset Statistics
@@
To construct MMGenBench-Domain, we randomly sampled $10,000$ images from the JourneyDB validation set. By utilizing the MMGenBench-pipeline, we can evaluate the image understanding and descriptive performance of LMMs within this domain, obviating the need for additional data.
@@ 3.3 Data Collection
To construct MMGenBench-Test, we first extract all images from the JourneyDB test set and process them using the process shown in Fig. 5. For the domain-specific dataset, we use the JourneyDB validation set for its construction. Subsequently, we will elaborate on each processing step.
@@
Pattern Extraction. We extract image patterns from existing images to measure the understanding and description ability of LMMs across various image categories. Given that we possess only images but no additional information, we utilize the powerful GPT-4o model to extract and analyze the underlying image patterns. The task is executed by GPT-4o in three sequential steps: 1) providing a detailed description of the image; 2) annotating the possible patterns based on the image features and description; and 3) explaining the rationale behind the annotated image patterns. By utilizing our carefully designed prompts (refer to Appendix B.3), GPT-4o can perform the task effectively. An example is presented in Fig. 10 in Appendix B.2, wherein the image description, alongside the extracted image patterns and rationales, aligns accurately with the image content. By annotating each image in the JourneyDB test set, we ultimately obtained a total of $1868$ image patterns.
@@
Annotation Summary. By leveraging GPT-4o, we extracted an extensive range of image patterns and observed that the semantic meanings across different patterns may be consistent. We conducted thorough statistics of the image patterns, by quantifying the occurrences of each type and ranking them in descending order (as seen in Appendix B.1). Subsequently, GPT-4-Turbo was utilized to generate the summary. To ensure the accuracy and validity of the extracted patterns, we carefully designed model prompts for this task (shown in Appendix B.3). This procedure ultimately resulted in a hand-crafted list of $13$ distinct image patterns, each thoroughly described in Appendix B.1.
@@
Image Re-annotation. We conducted a re-annotation of the JourneyDB test set using the previously identified $13$ image patterns. Similar to Image Pattern Extraction, GPT-4o was employed to generate descriptions for each image. Subsequently, these descriptions were annotated based on the predefined image patterns, accompanied by the underlying reasons. A key difference in this process is that re-annotation is restricted to the specified $13$ image patterns, excluding any other patterns. Given the extensive number of images in the test set, it is essential to select a subset of the images for subsequent operations. The re-annotated images were systematically classified into distinct image patterns, each containing a specified number of images. To construct the final dataset, we randomly sampled a subset of images from each pattern with a selection probability of $\frac{100}{N}$, where $N$ represents the total number of images within the pattern. To mitigate potential annotation errors, we manually verified and annotated each image. The final dataset was established by voting between the model predictions and annotations.
@@
Domain Data Collection. Constructing high-quality datasets is an extremely expensive task. Therefore, it is crucial to develop an efficient and convenient method for creating domain-specific datasets when evaluating the image understanding and description capabilities of LMMs in a new domain. By leveraging the MMGenBench-pipeline, our method enables a seamless evaluation of LMM performance across various domains, only depending on the availability of domain-specific images. In this study, we randomly select $10,000$ images from the JourneyDB validation set to create our domain-specific dataset. To quantify LMM performance within this domain, we calculate both FID-Score and SIM-Score.